OpenStudy (anonymous):

1.Solve S = 4v2 for v

OpenStudy (anonymous):

\[\frac{s}{4}=\frac{4v^2}{4}\] \[v^2=\frac{s}{4}\] \[v=\sqrt{\frac{s}{4}}\] \[v=\pm \frac{\sqrt{s}}{2}\]

OpenStudy (anonymous):

4.Solve V = π r2h for h